OpenStudy (anonymous):

just a simple true or false question please help the distance between points (x1 y1) and (4 8) is the square root of (x1 - 8)^2 + ( y 1 - 4)^2 the 1's by the x's and y's are lower then the variables btw

OpenStudy (kirbykirby):

\[ \sqrt{(x_1-8)^2+(y_1-4)^2} ~ ?\]

OpenStudy (anonymous):

I know the distance formula but how do I do it with only one set of end points?

OpenStudy (kirbykirby):

Well you can't know the distance with just one set of points. But, the question is asking if that formulation for the distance between \((x_1, y_1)\) and \((4, 8)\) is true?

OpenStudy (anonymous):

yeah

OpenStudy (kirbykirby):

False then. For it to be true, it would have to be \(\sqrt{(x_1-4)^2+(y_1-8)^2} ~ ?\) because the formula is \(\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} ~ ?\) (it doesnt't matter if you switch the order of x1 and x2 , or y1 and y2 to make it (x2 - x1) )

OpenStudy (anonymous):

Okay thank you

OpenStudy (kirbykirby):

:) yw